# Automatic Hardware Pragma Insertion in High-Level Synthesis: A Non-Linear Programming Approach

### v1

@@ 5. Non-Linear Formulation for Pragma Insertion
The II for each loop, the dependencies, the properties of the loops, the TC, the iteration latency of the parallel operations and the reduction operations and the number of DSPs per operation per statements are computed at compile time with PolyOpt-HLS (Pouchet et al., 2013a) and used as constants in the NLP problem.
@@ 5.2. Modeling Compiler Transformations
Pipeline Rules  Vitis HLS unrolls all loops under the pipelined loop. This implies that all loop $l$ under the pipelined loop must have $loop_{l}\_UF==TC_{l}$ (Eq. 15). Considering this constraint, it is important to note that only one pipelined loop can exist per nested loop (Eq. 5). If multiple pipelined loops were present, the loops beneath the first pipelined loop would be unrolled instead.
Memory Transfer Rules  Merlin automates the process of transferring data on-chip and applying array partitioning. The tool caches data on-chip and packs it in chunks of up to 512 bits, enabling efficient transfer speed. When the data is already present on-chip it can be reused provided that resource constraints are satisfied. The compiler caches on-chip the data only above the loop pipelined (Eq. 14).
Dependencies  Loop-carried dependencies are managed using constraints (Eq 8). If a loop has a dependency distance of $n$, this means that if we unroll with an unroll factor $uf>n$ this is equivalent to unrolling the loop with a factor $uf=n$ because the statements corresponding to the iteration $\{uf+1,\dots,n\}$ will be executed only after the first n statements are executed due to the dependency. Loop-independent data dependencies are managed at the objective function, as elaborated in Section 5.4.
Supplementary Rules  In addition, we add the constraints for the maximum unrolling (Eq. 10), the divisibility of the problem size of the unroll factors (Eq. 6), the tile size (Eq. 7) and the maximum array partitioning (Eq. 13). For this last, this corresponds to adding an upper bound to the product of the UF of all the loops which iterate the same array on different dimensions.
@@ 5.3. Constraints
(11) $DSPs\_used_{optimistic}=\sum_{op\in\{+,-,*,/\}}\max_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})$
@@ 5.5. Example
1Loop0: for(i=0; i<2100; i++) S0:y[i] = 0;
2Loop1: for(i=0; i<1900; i++) { S1:t[i] = 0;
3 Loop2: for(j=0; j<2100; j++) S2:t[i]+=A[i][j]*x[j];
4 Loop3: for(j=0; j<2100; j++) S3:y[j]+=A[i][j]*t[i];}
Listing 2: Atax Large
We now employ the Atax kernel as an illustration. $S0$ and $S3$ do not have inter-iteration dependencies within their respective loops, namely $Loop0$ and $Loop3$. Therefore, it is possible to pipeline $Loop0$ and $Loop3$ with an $II\geq 1$. In other words, for all iterations within these loops, there are no dependencies on previous iterations within the same loop. Loop1 and Loop2 are reduction loops, and the reduction operations is an addition which has an IL of $IL_{+}$ cycles. So the II ¿= $IL_{+}$. If loop1 is pipelined, there is a dependency between S1, S2 and S3 Between S1 and S2 the distance is 1, so we just add IL_S1 and IL_S2. Between S2 and S3 the distance is log(N) so the final equation will be IL_S1 + IL_S2 * log(N) + IL_S3 for the loop body cycle. If statements can be run at the same time (i.e., there is no dependency) it is a max instead of an addition. If we encounter code like $for(j=0;j<N;j++)y[j]=y[j-2]+3;$, a straightforward approach to handling this type of dependency is to impose a constraint such as $loop_{l}\_UF\leq 2$, which is represented by equation 8. In this case, due to dependencies an $UF>2$ is similar to $UF=2$.
@@ 6. Design Space Exploration
for $max\_array\_partitioning\in\{\infty,2048,1024,512,256,128,64,32,16,8,1\}$ do
@@ 7.1. Setup
We use kernels from Polybench/C 4.2.1 (PolyBench, [n. d.]). In addition we add a kernel of Convolution Neural Network (CNN). A single-precision floating point is utilized as the default data type in computations. Computations operate on medium and large datasets from PolyBench/C (PolyBench, [n. d.]) in order to have kernels with large footprint and have a large enough space to explore. The problem size and loop order of CNN are J,I=256, P,Q=5 H,W=224. A description of each benchmark can be found in Table 2. The ludcmp, deriche and nusinnov kernels are not present as PolyOpt-HLS (Pouchet et al., 2013b) does not handle negative loop stride. Cholesky and correlation contains a sqrt() operation which we do not support currently. Finally, we removed FDTD-2D because it exposed a bug in Merlin, and this generated a program where data dependencies are not fully preserved.
@@
We analyze the kernels and automatically generate each NLP problem with a version of PolyOpt-HLS (Pouchet et al., 2013a). We modified and extended for our work. Employing the AMPL description language to solve the NLP problems, we ran the commercial BARON solver (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017) version 21.1.13. For our experiments, we utilize 2 Intel(R) Xeon(R) CPU E5-2680 v4 @ 2.40GHz and 252GB DDR4 memory.
@@ 7.2. Experimental Evaluation
We compare our method with AutoDSE (Sohrabizadeh et al., 2021), described in Section 2, and we automatically generate the space of AutoDSE with the command ds_generator. We replace the UF and tile size by all the UF and tile size which divide the TC in order to have the same space. AutoDSE doesn’t impose any constraints on parallelism or the maximum array partitioning. It employs an incremental exploration approach, enabling it to make compiler-specific pragma selections.
Table 2 displays the space size of each design. The DSE is done in 4 parts with two threads for each (default parameter), with a timeout for the generation of the HLS report of 180 minutes, and a timeout of the DSE of 600 minutes (not always respected cf. Table 2). For our method we take the same parameters, and we add a timeout for BARON of 30 minutes.
@@ 7.3. Comparison with AutoDSE
Table 2. Comparison with AutoDSE
Table 2 shows the comparison with AutoDSE. NL, ND, S, and Space S are respectively the number of loops, the number of polyhedral dependencies (WaR, WaW, RaW), problem size (L for Large and M for Medium) and space size. For each method we compute the throughput (GF/s) in GFLOPs per second, the total time of the DSE (T) in minutes, the number of designs explored (DE) and the number of designs timeout (DT). In addition, for AutoDSE we add the number of design that are early rejected/prune (ER) as AutoDSE prunes the design when AMD/Xilinx Merlin can not apply one of the pragmas, due to its analysis limitations.
The performance of the kernel evaluated show significant improvements in both time and throughput. The time of the DSE is 5.69x faster on average (3.70x for geo-mean) and the throughput is 17.24x higher on average (2.38x for geo-mean) for the kernel evaluated. For almost all (46/47) kernels and problem sizes the method identifies a design with a throughput similar to (+/- 2%), or better than, AutoDSE. We have a slight slowdown for Doitgen Large because NLP-DSE explores the design found by the NLP with a maximum array partitioning of 2048 which timeouts, and then 1024 which is the best design found. However AutoDSE finds a design with a maximum array partitioning of 1280. By changing the maximum array partitioning to 1280 we find the same configuration as AutoDSE. Thus it is possible to obtain designs with a better performance but at the cost of a longer search. For all kernels and problem sizes, except Durbin, NLP-DSE is faster than AutoDSE. AutoDSE prunes all configurations of Durbin which explains the speed of AutoDSE for this kernel.
We can observe a difference of the performance for the same kernel in function of the problem size. If we take the example of 2mm, the difference has many factors. First as the footprint of the kernel becomes more important, it begins overusing the BRAMs. A large parallelism requires a bigger array partitioning which considerably increases the number of BRAMs and uses more BRAMs than available. Additionally, for large problems with high levels of parallelism, there are multiple instances of timeouts observed. Furthermore, the compilers applied the pragmas more efficiently for smaller problem sizes. We observe twice as many kernels where the pragmas are not applied as expected for the large problem size.
For Atax Large (Lst. 2), AutoDSE explores 166 designs of which 106 are early rejected and 30 timeout. AutoDSE starts by partially unrolling Loops 2 and 3 and will then attempt to do a coarse-grained parallelization on Loop 1 with all divisors, which is impossible due to dependencies. Although AutoDSE manages to prune/early reject the designs because Merlin cannot apply the pragmas, it still requires several minutes of compilation by Merlin for each unroll factor. In parallel, AutoDSE tries to pipeline the outermost loops (and therefore unroll the innermost loops) which creates numerous timeouts. Although the first two designs timeout due to too high level of parallelism, NLP-DSE allows us to find a configuration with the innermost loops unrolled with a UF=700. This allows us to find a design with a 3.46x higher throughput in 11.34x less time.
@@ 7.4. Accuracy
The tightness of the lower bound estimation relies on the correct application of pragma directives such as pipeline and parallel. It also assumes that Merlin can efficiently transfer memory from off-chip to on-chip using 512-bit chunks. Finally, it assumes that Merlin optimally handles the transfer of memory from off-chip to on-chip. Figures 1a and 1b compare the measured HLS latency for every synthesizable design explored during our DSE with its predicted latency per solving the NLP. The Y-axis represents log(latency) and the X-axis the rank of the design sorted by HLS latency. For Fig 1b we exclude designs when we detect that the pragma parallel and pipeline are not applied as defined by Merlin. We observe that about half of the designs have at least one pragma not applied, leading logically to a larger difference between measured and predicted latency. Generally speaking, for parallelization pragmas, Merlin is more restrictive for coarse-grained parallelization, in many cases these pragmas are not applied. We also observe certain cases where the partitioning is not done correctly which does not allow a pipeline with II=1 when it is theoretically possible. For Figure 1b we observe a better overall accuracy, albeit imperfect. These differences are due in large part to how Merlin eventually implemented memory transfers, which we model optimistically.
(a) when some pragmas are not applied
(b) when all pragmas are applied
Figure 1. Comparison of the HLS Latency vs. Lower Bound
@@
We observe in Fig. 1a and Fig. 1b one configuration where the lower bound property is not maintained (shown in red). This corresponds to a configuration of the Heat-3d kernel, where the pragma loop_flatten has been applied automatically, which changes the program structure. Overall unless Vitis applies loop_flatten automatically, which we do not model, our estimate is a lower bound for the cases evaluated.

### v2

@@ 5. Non-Linear Formulation for Pragma Insertion
The II for each loop, the dependencies, the properties of the loops, the TC, the iteration latency of the parallel operations and the reduction operations and the number of DSPs per operation per statements are computed at compile time with PolyOpt-HLS (Pouchet et al., 2013b) and used as constants in the NLP problem.
@@ 5.1. Variables
Set Description
$\mathcal{L}$ the set of loops
$\mathcal{A}$ the set of arrays
$\mathcal{S}$ the set of statements
$\mathcal{O}_{s}$ the list of operations of the statements $s$
$\mathcal{O}_{s_{par}}$ the operation which can be done in parallel, i.e., does not have any loop-carried dependence
$\mathcal{P}$ the set of different possible pipeline configurations
$\mathcal{L}^{p}_{pip}$ the set of loops pipelined
$\mathcal{L}^{p}_{under\_pip_{l}}$ the set of loops under a loop pipelined
$\mathcal{L}^{p}_{above\_pip_{l}}$ the set of loops above the loop pipelined $l$
$\mathcal{L}_{s}$ the set of nested loops which iterate the statement $s$
$\mathcal{C}_{a_{d}}$ the set of loops which iterates the array $a$ at the dimension $d$
$d_{l}$ the maximum dependency distance of the loop $l$
Variable Description
$loop_{l}\_UF$ Unroll factor of the loop $l$
$loop_{l}\_tile$ TC of the innermost loop after strip-mining of the loop $l$
$loop_{l}\_pip$ Boolean to know if the loop $l$ is pipelined
$loop_{l}\_cache\_array\_a$ Boolean to know the the array $a$ is transferred on-chip before the loop $l$
Constant Description
$TC_{l}$ Trip Count of the loop $l$
$II_{l}$ II of the loop $l$
$IL_{par}$ Iteration Latency of the operations whitout dependencies of the statement $s$
$IL_{red}$ Iteration Latency of the operations whit dependencies of the statement $s$
$DSP_{s_{op}}$ Number of DSP used for the statement $s$ for the operation $op$
$DSP_{available}$ Number of DSP available for the FPGA used
$MAX\_PARTITIONING$ Maximum array partitioning defined by the user or the DSE (cf. Section 6)
$footprint\_array_{a}\_loop_{l}$ Footprint of the array $a$ if transferred on-chip before the loop $l$
Table 4. Description of the sets, variables, and constants utilized in the formulation of the Nonlinear Problem (NLP) aimed at modeling latency and resource consumption of a design
Table 4, summarize the sets, variable and constants we use.
@@ 5.2. Modeling Compiler Transformations
Pipeline Rules
Vitis HLS unrolls all loops under the pipelined loop. This implies that all loop $l$ under the pipelined loop must have $loop_{l}\_UF==TC_{l}$ (Eq. 15). Considering this constraint, it is important to note that for each statement, only one of the loops that iterate the statement can have a pragma pipeline (Eq. 5). If multiple pipelined loops were present, the loops beneath the first pipelined loop would be unrolled instead.
Memory Transfer Rules
Merlin automates the process of transferring data on-chip and applying array partitioning. The tool caches data on-chip and packs it in chunks of up to 512 bits, enabling efficient transfer speed. When the data is already present on-chip it can be reused provided that resource constraints are satisfied. The compiler caches on-chip the data only above the loop pipelined (Eq. 14).
Dependencies
Loop-carried dependencies are managed using constraints (Eq 8). If a loop has a dependency distance of $n$, this means that if we unroll with an unroll factor $uf>n$ this is equivalent to unrolling the loop with a factor $uf=n$ because the statements corresponding to the iteration $\{uf+1,\dots,n\}$ will be executed only after the first n statements are executed due to the dependency. Loop-independent data dependencies are managed at the objective function, as elaborated in Section 5.4.
Supplementary Rules
In addition, we add the constraints for the maximum unrolling (Eq. 10), the divisibility of the problem size of the unroll factors (Eq. 6), the tile size (Eq. 7) and the maximum array partitioning (Eq. 13). For this last, this corresponds to adding an upper bound to the product of the UF of all the loops which iterate the same array on different dimensions.
@@ 5.3. Constraints
(11) $DSPs\_used_{optimistic}=\sum_{op\in\{+,-,*,/\}}\max_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})\leq DSP_{available}$
@@ 5.5. Example
1Loop0: for(i=0; i<2100; i++)
2 S0:y[i] = 0;
3Loop1: for(i=0; i<1900; i++) {
4 S1:t[i] = 0;
5 Loop2: for(j=0; j<2100; j++)
6 S2:t[i]+=A[i][j]*x[j];
7 Loop3: for(j=0; j<2100; j++)
8 S3:y[j]+=A[i][j]*t[i];
9}
Listing 10: AtAx code for Large problem size: $t=A*x;y=A*t$
We now employ the AtAx kernel as an illustration. $S0$ and $S3$ do not have inter-iteration dependencies within their respective loops, namely $Loop0$ and $Loop3$. Therefore, it is possible to pipeline $Loop0$ and $Loop3$ with an $II\geq 1$. In other words, for all iterations within these loops, there are no dependencies on previous iterations within the same loop. Loop1 and Loop2 are reduction loops, and the reduction operation is an addition which has an IL of $IL_{+}$ cycles. So the II ¿= $IL_{+}$. If loop1 is pipelined, there is a dependency between S1, S2 and S3 Between S1 and S2 the distance is 1, so we just add IL_S1 and IL_S2. Between S2 and S3 the distance is log(N) so the final equation will be IL_S1 + IL_S2 * log(N) + IL_S3 for the loop body cycle. If statements can be run at the same time (i.e., there is no dependency) it is a max instead of an addition. If we encounter code like $for(j=2;j<N;j++)y[j]=y[j-2]+3;$, a straightforward approach to handling this type of dependency is to impose a constraint such as $loop_{l}\_UF\leq 2$, which is represented by equation 8. In this case, due to dependencies an $UF>2$ is similar to $UF=2$.
@@ 6. Design Space Exploration
for $max\_array\_partitioning\in Space\_Array\_Partitioning$ do
@@ 7.1. Setup
We use kernels from Polybench/C 4.2.1 (PolyBench, [n. d.]). The complexities and sizes of the problems are detailed in Table 8 located in the Appendix. In addition we add a kernel of Convolution Neural Network (CNN). A single-precision floating point is utilized as the default data type in computations to compare to AutoDSE. Computations operate on medium and large datasets from PolyBench/C (PolyBench, [n. d.]) in order to have kernels with large footprint and have a large enough space to explore. Selecting medium and large problem sizes is crucial to accurately reflect the complexities encountered in various fields such as scientific simulations, data analytics, and artificial intelligence. These sizes pose challenges that mirror the practical limitations of memory transfer, where efficiently managing data movement becomes paramount due to its potential to bottleneck performance. In such contexts, the careful orchestration of memory transfers is essential to ensure optimal resource utilization and prevent computational inefficiencies. Additionally, the scale of these problems often exceeds on-chip memory capacity, necessitating strategies for effectively handling data footprints that surpass available memory, further emphasizing the need for meticulous memory management and optimization techniques. The problem size and loop order of CNN are J,I=256, P,Q=5 H,W=224. A description of each benchmark can be found in Tables 8 and 5. The ludcmp, deriche and nusinnov kernels are not present as PolyOpt-HLS (Pouchet et al., 2013a) does not handle negative loop stride. Cholesky and correlation contains a sqrt() operation which we do not support currently. Finally, we removed FDTD-2D because it exposed a bug in Merlin, and this generated a program where data dependencies are not fully preserved.
A double-precision floating point is utilized as the default data type in computations to compare to HARP. We chose to use the problem size used by HARP in order to reuse their model. The problem size and kernel use by HARP can be found on the Table 9.
@@
We analyze the kernels and automatically generate each NLP problem with a version of PolyOpt-HLS (Pouchet et al., 2013b). We modified and extended for our work. Employing the AMPL description language to solve the NLP problems, we ran the commercial BARON solver (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017) version 21.1.13. For our experiments, we utilize 2 Intel(R) Xeon(R) CPU E5-2680 v4 @ 2.40GHz and 252GB DDR4 memory.
@@ 7.2. Experimental Evaluation
7.2.1. AutoDSE
We compare our method with AutoDSE (Sohrabizadeh et al., 2021), described in Section 2, and we automatically generate the space of AutoDSE with the command ds_generator. We replace the UF and tile size by all the UF and tile size which divide the TC in order to have the same space. AutoDSE does not impose any constraints on parallelism or the maximum array partitioning. It employs an incremental exploration approach, enabling it to make compiler-specific pragma selections.
Table 5 displays the space size of each design. The DSE is done in 4 parts with two threads for each (default parameter), with a timeout for the generation of the HLS report of 180 minutes, and a timeout of the DSE of 600 minutes (not always respected cf. Table 5). For our method we take the same parameters, and we add a timeout for BARON of 30 minutes. The space given as input to NLP-DSE is $\{\infty,2048,1024,512,256,128,64,32,16,8,1\}$.
7.2.2. HARP
The evaluation vs. HARP is done with the same parameters as the evaluation vs. AutoDSE. We change the space given as input due to the small problem size and we choose $\{\infty,1024,750,512,256,128,64,32,16,8,1\}$.
We run HARP for one hour in order to have a similar DSE time as NLP-DSE . This enables the exploration of an average of 75,000 distinct pragma configurations for each kernel. HARP’s DSE method navigates the space by iteratively adjusting the pragma in a bottom-up manner. It synthesizes the top 10 designs discovered by the DSE, employing a timeout of 3 hours for the HLS compiler, similar to the approach used in the NLP-DSE framework.
@@ 7.3. Comparison with AutoDSE
Figures 2 and 3 show the comparison with AutoDSE for Large and Medium problem size respectively. Table 5 shows the details of the comparison with AutoDSE. NL, ND, S, and Space S are respectively the number of loops, the number of polyhedral dependencies (WaR, WaW, RaW), problem size (L for Large and M for Medium) and space size. For each method we compute the throughput (GF/s) in GFLOPs per second, the total time of the DSE (T) in minutes, the number of designs explored (DE) and the number of designs timeout (DT). In addition, for AutoDSE we add the number of design that are early rejected/prune (ER) as AutoDSE prunes the design when AMD/Xilinx Merlin can not apply one of the pragmas, due to its analysis limitations.
The performance of the kernel evaluated show significant improvements in both time and throughput. The time of the DSE is 5.69x faster on average (3.70x for geo-mean) and the throughput is 17.24x higher on average (2.38x for geo-mean) for the kernel evaluated. For almost all (46/47) kernels and problem sizes the method identifies a design with a throughput similar to (+/- 2%), or better than, AutoDSE. We have a slight slowdown for Doitgen Large because NLP-DSE explores the design found by the NLP with a maximum array partitioning of 2048 which timeouts, and then 1024 which is the best design found.
(a) GF/s
(b) DSE Time (min)
Figure 2. Comparison between the throughput (GF/s) and Design Space Exploration (DSE) time (min) of NLP-DSE and AutoDSE for large problem sizes in Polybench.
{adjustwidth}
-1in-1in
@@
Table 5. Comparison of DSE time and Throughput for NLP-DSE , NLP-DSE-FS, and AutoDSE across Polybench kernels for different problem sizes
(a) GF/s
(b) DSE Time (min)
Figure 3. Comparison between the throughput (GF/s) and Design Space Exploration (DSE) time (min) of NLP-DSE and AutoDSE for medium problem sizes in Polybench.
However AutoDSE finds a design with a maximum array partitioning of 1280. By changing the maximum array partitioning to 1280 we find the same configuration as AutoDSE. Thus it is possible to obtain designs with a better performance but at the cost of a longer search. For all kernels and problem sizes, except Durbin, NLP-DSE is faster than AutoDSE. AutoDSE prunes all configurations of Durbin which explains the speed of AutoDSE for this kernel.
We can observe a difference of the performance for the same kernel in function of the problem size. If we take the examples of 2mm and 3mm, the difference has many factors. First as the footprint of the kernel becomes more important, it begins overusing the BRAMs. A large parallelism requires a bigger array partitioning which considerably increases the number of BRAMs and uses more BRAMs than available. Additionally, for large problems with high levels of parallelism, there are multiple instances of timeouts observed. Furthermore, the compilers applied the pragmas more efficiently for smaller problem sizes. We observe twice as many kernels where the pragmas are not applied as expected for the large problem size.
For AtAx Large (Lst. 10), AutoDSE explores 166 designs of which 106 are early rejected and 30 timeout. AutoDSE starts by partially unrolling Loops 2 and 3 and will then attempt to do a coarse-grained parallelization on Loop 1 with all divisors, which is impossible due to dependencies. Although AutoDSE manages to prune/early reject the designs because Merlin cannot apply the pragmas, it still requires several minutes of compilation by Merlin for each unroll factor. In parallel, AutoDSE tries to pipeline the outermost loops (and therefore unroll the innermost loops) which creates numerous timeouts. Although the first two designs timeout due to too high level of parallelism, NLP-DSE allows us to find a configuration with the innermost loops unrolled with a UF=700. This allows us to find a design with a 3.46x higher throughput in 11.34x less time.
@@
7.4. Comparison with HARP
Utilizing the PolyBench problem size of HARP allows for direct reusability of the model, facilitating comparison and benchmarking against the framework HARP. This also allows for the utilization of data that will enable achieving the best results with HARP.
Evaluating on other problem sizes would have required the creation of a database to at least fine-tune the model with the kernel and problem size in question.
Figure 4 shows the comparison with HARP for Small and Medium problem size. Table 9 in Appendix, shows the details of the comparison.
The throughput is 1.45x higher on average (1.20x for geo-mean) for the kernel evaluated in similar DSE time. For 20/23 kernels the method identifies a design with a throughput similar to (+/- 10%), or better than, HARP.
We note a variation in the enhancement of performance compared to the evaluation in Section 7.3 vs. AutoDSE, stemming from various factors. Primarily, the breadth of the exploration space plays a significant role. HARP has the capacity to traverse an average of 150,000 designs, enabling it to nearly exhaustively explore the entire space. Additionally, HARP is trained and/or fine-tuned with precise knowledge of the kernel and problem size, granting it deep insight into scenarios where pragmas are not applied and have enough training on these specific kernel to estimate the latency. This confers an advantage over AutoDSE, which treats the compiler as a black box.
Figure 4. Comparison between the throughput (GF/s) of NLP-DSE and HARP for small and medium problem sizes in Polybench.
In the case of the medium-sized mvt kernel, a notable slowdown occurs because Merlin is transferring the array A twice. This array comprises 159,900 double-precision floats. The kernel is already constrained by memory bandwidth when transferring A once. The duration for a single transfer of A is 20,000 cycles, utilizing a bitwidth of 512 (maximum), amounting to a total of 40,000 cycles. Consequently, the kernel’s overall latency is 40,726 cycles. However, HARP successfully identified a configuration that enables Merlin to transfer A only once.
With Gemver, we can attain a speedup by leveraging our capability to explore the entire space within a single optimization problem. The space of Gemver with medium size encompasses over $10^{11}$ designs, making it impractical to thoroughly explore, even with HARP’s estimation per design hovering around the millisecond range.
@@ 7.5. Accuracy
The tightness of the lower bound estimation relies on the correct application of pragma directives such as pipeline and parallel. It also assumes that Merlin can efficiently transfer memory from off-chip to on-chip using 512-bit chunks. Finally, it assumes that Merlin optimally handles the transfer of memory from off-chip to on-chip. Figures 5a and 5b compare the measured HLS latency for every synthesizable design explored during our DSE with its predicted latency per solving the NLP. The Y-axis represents log(latency) and the X-axis the rank of the design sorted by HLS latency. For Fig 5b we exclude designs when we detect that the pragma parallel and pipeline are not applied as defined by Merlin. We observe that about half of the designs have at least one pragma not applied, leading logically to a larger difference between measured and predicted latency. Generally speaking, for parallelization pragmas, Merlin is more restrictive for coarse-grained parallelization, in many cases these pragmas are not applied. Coarse-grained pragmas are typically not applied to kernels that do not have an outermost reduction loop and thus can theoretically have coarse-grained parallelization, which is present in most linear algebra kernels such as 2mm, 3mm, gemver, etc. We also observe certain cases where the partitioning is not done correctly which does not allow a pipeline with II=1 when it is theoretically possible. For Figure 5b we observe a better overall accuracy, albeit imperfect. These differences are due in large part to how Merlin eventually implemented memory transfers, which we model optimistically. Internally, Merlin transforms the size of the arrays according to the program’s unroll factors and in certain cases does not allow transfers with a bitwidth of 512 bits.
(a) Comparison of the latency between the design reported in the HLS report and the lower bound estimate provided by the nonlinear problem for all explored designs
(b) Comparison of the latency between the design reported in the HLS report and the lower bound estimate provided by the nonlinear problem specifically for cases where pragmas were applied
Figure 5. Comparison of the latency between the design reported in the HLS report and the lower bound estimate provided by the nonlinear problem for all explored designs and specifically for cases where pragmas were applied.
@@
We observe in Fig. 5a and Fig. 5b one configuration where the lower bound property is not maintained (shown in red). This corresponds to a configuration of the Heat-3d kernel, where the pragma loop_flatten has been applied automatically, which changes the program structure. Overall unless Vitis applies loop_flatten automatically, which we do not model, our estimate is a lower bound for the cases evaluated. Our model can easily implement the automatic flattened loop optimization: We must multiply the TC of the loop pipeline by the TC of all the perfectly nesteed loops above pipeline loop (and remove them in the first products). Because this optimization is rarely applied, we prioritize having a tight lower bound.
Additionally, we evaluate the number of DSE steps needed to achieve the design with the best Quality of Results (QoR) of our DSE and the number of syntheses required before terminating the DSE due to finding a lower bound (LB) greater than the latency of an already synthesized kernel, sorting by latency estimation provided by the NLP in ascending order.
Kernel To find the best QoR To find a LB ¿ than HLS result
Problem Size Large Medium Large Medium
2mm 4 6 13 9
3mm 7 2 11 7
atAx 6 2 18 14
bicg 2 2 16 18
covariance 12 8 16 18
doitgen 1 0 4 10
durbin 0 16 21 20
fdtd-2d 12 1 18 8
floyd-warshall 8 16 20 20
gemm 7 4 13 7
gemver 5 9 12 11
gesummv 0 4 14 16
gramschmidt 8 14 10 16
heat-3d 17 20 19 15
jacobi-1d 16 0 18 10
jacobi-2d 6 18 16 20
lu 18 0 20 20
mvt 1 0 4 6
seidel-2d 17 10 20 20
symm 12 10 16 21
syr2k 13 18 16 20
syrk 17 17 18 20
trisolv 0 0 17 20
trmm 16 4 20 17
Table 6. The count of designs evaluated to identify the HLS design yielding the optimal Quality of Results (QoR), and the count at which the Design Space Exploration (DSE) ceases upon discovering a lower bound (LB) surpassing the latency of a design already synthesized with the HLS compiler
The results are presented in Table 6. On average, it takes 8 steps of the DSE to discover the design with the best QoR and 15 steps to terminate the DSE. We can observe that for some kernel we find the design with the best QoR at the first iteration of the DSE (which correspond to first shoot method in Table 5) but the DSE need more stop to stop the guarantee that we cannot obtain better latency, which implies that the lower bound is not perfectly tight.